# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

### v2

@@ 3.1 Case-Enhanced Law Article Knowledge Graph Construction
Figure 2: Schema of CLAKG. Ellipses of different colors represent
different entity types.
Figure 3: Example of CLAKG.
@@
Key information Key information in law article extracted by large language model
@@
Agree With A case agree with several pieces of key information
Applicable Law The law mentioned in the judgment documents that matches the case
@@
Reason Case Documents contain the reason of lawsuit in the case
Detail The name of a case corresponds to the specific information of the case
As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Case Knowledge Graph (ACKG). The LAKG includes original article of law, law article id, and key information. The key information is extracted from the original article of law using a LLM. The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case. The description of the case process is derived from the content of the judgment documents processed using a LLM.
Based on the reference law articles written in the judgment documents, it is possible to link the name of a case node with law article id nodes. We analyze the correlation between case information and the corresponding key information of applicable law articles using a LLM and ultimately select up to five of the most relevant keywords for connection. To ensure the consistency and accuracy of the knowledge graph, we ignore any keywords that do not exist in the graph output by the large language model. Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
@@ 3.2 LLM and CLAKG Based Law Article Recommendation Framework
Figure 4: The LLM and CLAKG based law article recommendation framework.
The LLM and CLAKG based law article recommendation framework is shown in Figure 4. Before the start of the law article recommendation task, the CLAKG constructed in Section 3.1 needs to undergo graph embedding preprocessing using the Relational Graph Convolutional Network (RGCN) model mentioned in Section 3.2.1. The process of law article recommendation begins with the user inputting a newly emerged case. In Section 3.2.2, a LLM is used to match the k most relevant keywords to this new case (with k=8 as selected in this paper). Based on this, in Section 3.2.3, a candidate law article retrieval algorithm is employed to retrieve q candidate law articles that may be suitable for the new case (with q=5 as chosen in this paper). In Section 3.2.4, the information from the new case input by the user is integrated with the candidate law articles and historical case information obtained in Section 3.2.3 into a prompt, which is then passed to the LLM to complete the law article recommendation task.
@@ 3.2.1 Graph Embedding on CLAKG with RGCN
Graph embedding involves utilizing the topological structure information in knowledge graphs to assign embedding vectors to each node within the graph. The Graph Convolutional Network (GCN) model, initially proposed by [36], was the first graph neural network used for graph embedding. [37] improved upon the GCN, introducing the Relation Graph Convolution Neural Networks (RGCN), which assigns different weights to different relationships, enhancing the quality of graph embeddings.
@@
where $h_{i}^{(l)}$ represents the embedding vector of node $i$ in the $l$-th layer, $N_{i}^{r}$ represents all neighboring nodes of node $i$ that have relationship $r$. $c_{i,r}$ is a hyperparameter, which in this paper is taken as $\left|N_{i}^{r}\right|$. Both $W_{r}^{(l)}$ and $W_{0}^{(l)}$ are parameter matrices.
@@
The training process of the model is shown in Figure 5. Firstly, the system randomly initializes the embedding vectors for all nodes and relations. Based on this, the system updates the node embeddings using Equation 1. The system then calculates the score based on the DisMult decomposition, as described in Equation 2. This paper selects the cross-entropy function as the loss function, as described in Equation 3. When the triple is a positive example, meaning the information truly exists in the knowledge graph, the value of $y$ is 1. When the triple is a negative example, meaning the information does not exist in the graph, the value of $y$ is 0 .
$f(s,r,o)=e_{s}^{T}R_{r}e_{o}$ (2)
where $e_{s}$ and $e_{o}$ represent the embedding vectors of the nodes s and o, and $R_{r}$ is the embedding vector of the relation r.
$\operatorname{Loss}=$ $-\frac{1}{(1+\omega)|\hat{\varepsilon}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Bigg[y\log l(f(s,r,o))$ $+(1-y)\log\left(1-l(f(s,r,o))\right)\Bigg]$ (3)
where $\mathcal{J}$ represents all the datasets, and $l$ is the logistic activation function, as described in Equation 4.
$l(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}$ (4)
@@ 3.2.2 Matching Key Information Nodes in CLAKG with LLM
In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users. To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs. The prompts for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
@@
Output Example Public and private property; multiple thefts; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
@@ 3.2.3 Law Article and Case Retrieval Algorithm
1:key information node, the embedded vector dictionary obtained in Section 3.2.1
2:5 Original articles of law node
1:Initialize: law_article_distances = [] , law_article_set = {} $\triangleright$ Initializes a list to store distances and corresponding law articles and a set to store candidate law articles
2:for $e\in\text{key information node}$ do
3:  Case_set.add_node($x\mid x-[\text{key}]\rightarrow e$) $\triangleright$ $x$ is ‘Original article of penal law’ node
4:end for
5:for $\text{law\_article}\in\text{law\_article\_set}$ do
6:  $\text{distance}=0$ $\triangleright$ Initializes the distance score
7:  for $e\in\text{key information node}$ do
8:   $\text{distance}\mathrel{+}=\operatorname{dis}(\text{emb}[\text{law\_article}],\text{emb}[\text{e}])$ $\triangleright$ Calculate the distance score
9:    where $\operatorname{dis}(a,b)=\frac{a\cdot b}{\|a\|\times\|b\|}$ $\triangleright$ Cosine similarity formula
10:  end for
[…]
12:end for
[…]
14:return law_article_distances[:5] $\triangleright$ Output the top five candidate law articles
@@
The proposed method for retrieving candidate law article involves traversing all ‘key information’ nodes obtained in 3.2.2 and retrieving the ‘Original article of law’ nodes that are connected to the ‘key information’ nodes via ‘key’ relationships. These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article. Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis. To this end, each node in the container is evaluated by calculating the distance scores (cosine similarity) between the ‘Original article of penal law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
These distance scores are summed to yield a cumulative distance score between each ‘Original article of law’ node and the newly reported case. Ultimately, the top five ‘Original article of law’ nodes with the highest cumulative distance scores are selected as the candidate law articles, serving as reference points for the LLM in completing the law article recommendation task.
@@ 3.2.4 LLM-Enhanced Law Article Recommendation
Our method intends to incorporate reference case information and the original text of candidate law articles to eliminate the hallucinations of LLMs. After retrieving candidate law articles and relevant historical cases with the method described in Section 3.2.3, The knowledge graph is then queried for key details, and this integrated information helps determine the applicability of the candidate articles. The prompt used for the LLM in this chapter consists of expert roles, task descriptions, information on new criminal cases, reference case information, candidate law articles, and output examples.
After the user describes the case information to the system, the system returns the results generated by the LLM. The user can further inquire about the results provided by the LLM.
@@ 4.1 Datasets
The method proposed in this paper can be applied to any type of law articles in any country. For research convenience, the law articles datasets employed in this section is derived from the Amendment (XI) to the Criminal Law of the People’s Republic of China[38], ratified during the 24th session of the Standing Committee of the 13th National People’s Congress. This amendment is systematically structured into two sections: the General Provisions and the Specific Provisions, encompassing a total of 452 articles—101 in the General law articles and 351 in the Specific law articles. The CLAKG was constructed utilizing the methodology in Section 3.1.
The criminal case data analyzed in this study were obtained from the China Judgments Online [39] platform. Figure 6 presents one of the cases. In light of the fact that the Amendment (XI) to the Criminal Law of the People’s Republic of China came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date. During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited. As a result, certain law articles corresponded with only a limited number of cases, while some provisions had no corresponding cases at all. We curated a dataset comprising 1,170 case records, which collectively address law articles from 85 articles within the Specific Provisions. To optimize the dataset, we divided it in a manner that maximized the diversity of law articles represented in the test set. Specifically, 997 case records were designated for the training set and 173 for the test set. The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zoro shot number = 36). The Adjudicated Case Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.

### v3

@@ 3.1 Case-Enhanced Law Article Knowledge Graph Construction
Figure 2: Schema of CLAKG. Ellipses of different colors represent different entity types.
Figure 3: Example subgraph of CLAKG instantiated according to the schema defined in Figure 2.
@@
Key information Key information in law article
@@
Agree With A case agrees with several pieces of key information
Applicable Law The law that matches the case
@@
Reason Links a case to its prosecution reason stated in the judgment
Detail Links a case to its fact description
Figure 2 illustrates the schema of CLAKG, defining the entity types and relation types used in our system. An example subgraph instantiated according to this schema is shown in Figure 3. Tables 1 and 2 provide the details of all types of nodes and relationships in the CLAKG.
As shown in Figure 2, the Case-Enhanced Law Article Knowledge Graph (CLAKG) is composed of the Law Article Knowledge Graph (LAKG) and the Adjudicated Cases Knowledge Graph (ACKG). The LAKG includes original article of law, law article id, and key information. The key information is first extracted from the original article of law using an LLM, and then reviewed and, if necessary, refined by legal experts before being stored as nodes in CLAKG. The ACKG contains the name of a case, time of the court session, prosecution reason, and the specifics of a case. The description of the case process is derived from the content of the judgment documents processed using an LLM, and again experts check and adjust the extracted elements so that the resulting case nodes and their attributes faithfully reflect the original judgments.
Based on the reference law articles cited in the judgment documents, we link each name of a case node in ACKG to the corresponding law article id nodes. Furthermore, we analyze the correlation between the factual description of a case and the key information of the applicable law articles using an LLM, and select up to five of the most relevant key-information nodes for each case. Concretely, the LLM is asked to choose several items from the existing list of key-information strings for the cited articles; its answer is parsed as a semicolon-separated list and each substring is matched against this list. Any substring that does not exactly match an existing key-information node is discarded, and again no new nodes are added. After this automatic step, legal experts can confirm the suggested case–article links and key-information associations, correct inaccurate links, and add missing but legally important relations if the LLM has overlooked them. This human–machine collaborative filtering strategy ensures that the constructed CLAKG remains consistent with the predefined schema: noisy LLM outputs may reduce recall before expert correction, but they cannot introduce spurious concepts or inconsistent node types into the graph. Finally, each name of a case node is linked to several key information nodes.
@@ 3.2 LLM and CLAKG Based Law Article Recommendation Framework
Figure 4: The LLM and CLAKG based law article recommendation framework. The candidate law articles and historical cases retrieved from CLAKG are used to constrain the LLM’s reasoning and thereby mitigate hallucinations.
The LLM and CLAKG based law article recommendation framework is shown in Figure 4 and step 2 of Figure 1. Before the start of the law article recommendation task, the CLAKG constructed in Section 3.1 needs to undergo graph embedding preprocessing using the Relational Graph Convolutional Network (RGCN) model mentioned in Section 3.2.1. The process of law article recommendation begins with the user inputting a newly emerged case. In Section 3.2.2, a LLM is used to match the K most relevant keywords to this new case (with K=8 as selected in this paper). Based on this, in Section 3.2.3, a candidate law article retrieval algorithm is employed to retrieve q candidate law articles that may be suitable for the new case (with q=5 as selected in this paper). In Section 3.2.4, the information from the new case input by the user is integrated with the candidate law articles and historical case information obtained in Section 3.2.3 into a prompt, which is then passed to the LLM to complete the law article recommendation task, outputting at most $t$ law article IDs (with $t=1$ as selected in this paper).
@@ 3.2.1 Graph Embedding on CLAKG with RGCN
Given the CLAKG constructed in Section 3.1, we now describe how to compute node embeddings on this graph. Let $G=(V,E,R)$ denote the CLAKG, where each node $v\in V$ belongs to one of the entity types listed in Table 1 (original article of law, law article id, key information, case name, time of the court session, prosecution reason, and case details). Edges $e\in E$ are labeled by relation types $r\in R$ from Table 2, including Key (linking an article to its key information), Id (linking an article to its law article id), Applicable Law (linking a case to the articles cited in the judgment), Agree With (linking a case to the key information it matches), Occur in Time (linking a case to its decision date), Reason (linking a case to the reason of lawsuit), and Detail (linking a case to its detailed fact description). Together, these nodes and relations encode which law articles apply to which cases and through which key information, and they constitute the structural basis on which our graph embedding module operates.
The purpose of the graph embedding module in our framework is to exploit the heterogeneous, multi-relational structure of CLAKG and map every node $v\in V$ to a low-dimensional vector representation. These embeddings provide the matching capability in our framework: in Section 3.2, they are used to measure the similarity between the key-information nodes selected from a new case and all law-article nodes when ranking candidate law articles. To learn such representations on CLAKG, we adopt Relational Graph Convolutional Networks (RGCN) [36], which extend standard Graph Convolutional Networks (GCNs) [27] to multi-relational graphs by assigning different transformation matrices to different relation types and aggregating information from neighbors according to the graph structure.
@@
where $h_{i}^{(l)}$ represents the embedding vector of node $i$ in the $l$-th layer, $N_{i}^{r}$ represents all neighboring nodes of node $i$ that have relationship $r$. $c_{i,r}$ is a normalization constant, which in this paper is taken as $\left|N_{i}^{r}\right|$. Both $W_{r}^{(l)}$ and $W_{0}^{(l)}$ are parameter matrices.
@@
The training process is illustrated in Fig. 5. We first randomly initialize trainable embedding vectors for all entities and relations in the knowledge graph. Node representations are then updated by stacking $L$ layers of relational graph convolution (RGCN) as in Eq. 1, where each layer aggregates messages from relation-specific neighbors with appropriate normalization (e.g., by the neighborhood size) and applies a non-linear activation.
Given the final node embeddings, we compute the plausibility score of each triple $(s,r,o)$ using the DistMult [44] factorization:
$f(s,r,o)=\langle e_{s},e_{r},e_{o}\rangle=\sum_{k=1}^{d}e_{s,k}\,e_{r,k}\,e_{o,k},$ (2)
where $e_{s},e_{o}\in\mathbb{R}^{d}$ denote the embeddings of the subject node $s$ and object node $o$, respectively, and $e_{r}\in\mathbb{R}^{d}$ is the embedding of relation $r$.
We optimize the model using the binary cross-entropy (BCE) loss in Eq. 3. Let $\mathcal{T}$ denote the set of training examples consisting of observed positive triples and negatively sampled (corrupted) triples generated by replacing the head or tail entity with a randomly selected entity (i.e., head/tail corruption). The binary label $y\in\{0,1\}$ indicates whether a triple is positive ($y=1$) or a negative sample ($y=0$):
$\mathcal{L}=-\frac{1}{|\mathcal{T}|}\sum_{(s,r,o,y)\in\mathcal{T}}\Big[y\log\sigma(f(s,r,o))+(1-y)\log\!\big(1-\sigma(f(s,r,o))\big)\Big],$ (3)
where $\sigma(\cdot)$ is the logistic function defined in Eq. 4.
$\sigma(x)=\frac{1}{1+e^{-x}}=\frac{e^{x}}{e^{x}+1}.$ (4)
@@ 3.2.2 Matching Key Information Nodes in CLAKG with LLM
Output Example Public and private property; sale; committing a crime; causing damage; seriously disrupting public order; multiple thefts; lawfully performing duties
In this section, we aim to match the relevant key information nodes within CLAKG based on current case information provided by users. To facilitate this process, we have designed prompts for the LLM, which include the expert role, task description, details of the newly reported case, known key information nodes, and example outputs. The prompt template for LLM in the task of matching key information nodes in the CLAKG related to new case information is shown in Table 3.
During inference, we reuse the global list of key-information strings stored in CLAKG and present it to the LLM as the option set in the prompt. The LLM is instructed to output several items from this list, separated by semicolons. The output is post-processed by splitting on semicolons and performing exact matching against the list; substrings that do not match any existing key-information node are discarded and do not create new nodes.
@@ 3.2.3 Law Article and Case Retrieval Algorithm
1:Query key-information nodes $K$; embedding lookup table $\mathrm{emb}[\cdot]$ (Section 3.2.1); relation type $\{\textit{Key}\}$ in CLAKG
2:Top-q candidate $\{\textit{Original article of law}\}$ nodes (optionally with scores)
3:Initialize: $\leftarrow\emptyset; []\mathcal{L}\leftarrow[\ ] \triangleright SS$: candidate law-article set, $\mathcal{L}$: list of (score, law_article)
4:for $e\in K$ do
5:  $S\leftarrow S\cup\{\,x\mid(x\xrightarrow{\{\textit{Key}\}}e)\in E\ \wedge\ x\in\{\textit{Original article of law}\}\,\}$
6:end for
7:for $a\in S$ do
8:  $s\leftarrow 0$
9:  for $e\in K$ do
10:   $s\leftarrow s+\cos(\mathrm{emb}[a],\mathrm{emb}[e])$ $\triangleright$ $\cos(u,v)=\frac{u\cdot v}{\|u\|\|v\|}$
11:  end for
12:  $\mathcal{L}.\texttt{append}((s,a))$
13:end for
14:sort($\mathcal{L}$) by $s$ in descending order
15:return first q law articles in $\mathcal{L}$ (or their IDs)
@@
The proposed method for retrieving candidate law article involves traversing all ‘Key information’ nodes obtained in Section 3.2.2 and retrieving the ‘Original Article of Law’ nodes that are connected to the ‘Key Information’ nodes via ‘Key’ relationships. These nodes are then aggregated into a set container, with all ‘Original article of law’ nodes in the set representing the candidate law article. Given that the number of candidate law articles in the set container may be substantial, it is essential to select those law articles most analogous to the newly reported case for further analysis. To this end, each node in the container is evaluated by calculating the similarity scores (cosine similarity) between the ‘Original article of law nodes’ embedding vector and the multiple ‘key information’ nodes embedding vector obtained in 3.2.1.
These similarity scores are summed to yield a cumulative similarity score between each ‘Original Article of Law’ node and the newly reported case. Ultimately, the top five ‘Original Article of Law’ nodes with the highest cumulative similarity scores are selected as the candidate law articles, serving as reference points for the LLM in completing the law article recommendation task.
@@ 3.2.4 LLM-Enhanced Law Article Recommendation
As illustrated in Figure 6, once the CLAKG-based retrieval stages in Sections 3.2.2 and 3.2.3 have been completed, the system has obtained three types of structured evidence for a new case: (i) the textual description of the new case provided by the user, (ii) a small set of candidate law articles retrieved from CLAKG, and (iii) a set of reference historical cases linked to these candidate law articles. In this subsection we describe how this information is integrated into a Large Language Model so as to perform law article recommendation while mitigating hallucinations.
Figure 6: Flowchart of LLM-enhanced law article recommendation grounded in CLAKG. The LLM receives KG-retrieved candidate law articles and linked historical cases as verifiable context, which constrains its reasoning and helps mitigate LLM hallucinations.
To explicitly leverage the retrieved candidate law articles and reference historical cases as constraints for mitigating hallucinations, we design a structured prompt for the LLM-enhanced recommendation stage. For each new case, the prompt specifies the expert role, task description, the factual description of the case, the retrieved candidate law articles and their matched key-information nodes, as well as the reference historical cases linked to these articles. In addition, we impose strict output constraints that restrict the model to selecting from the candidate set and require it to explicitly state that no suitable law article is found rather than fabricating new ones. The detailed prompt components for CLAKG-grounded law article recommendation are summarized in Table 4.
Table 4: Prompts for LLM in the task of CLAKG-grounded law article recommendation.
Component Details
Expert Role Criminal-law expert specializing in applying Chinese criminal law articles to concrete cases.
Task Description Given the new case description, the candidate law articles, and the reference historical cases, recommend the law articles that are most applicable to the case. You must choose only from the candidate law articles and must not fabricate non-existent article numbers or law articles.
New Case Description Factual description of the new case provided by the user.
Candidate Law Articles For each candidate article: its ID and original article of law.
Reference Historical Cases Concise summaries of historical cases linked to each candidate article in CLAKG (the specifics of the case, prosecution reason, and judgment result).
Output Constraints Output at most $t$ law article IDs from the candidate set, each with a brief justification. If none of the candidates is suitable, explicitly state that no applicable law article is found, rather than inventing a new one.
Given this prompt, the LLM compares the new case with each candidate law article and its linked historical cases, and then produces a ranked list of recommended law articles together with short textual explanations. The law articles remaining after these checks are returned to the user as the Final Recommended Law Articles in Figure 6. The user can review these recommendations, request further explanations from the LLM (e.g., why a particular article applies or why similar articles are not chosen).
3.3 Closed-loop Human–Machine Collaboration and CLAKG Renewal
Algorithm 2 CLAKG Update with Human Feedback
1:CLAKG $G=(V,E,R)$, new case $c$, system law-article recommendations $A^{\text{sys}}$, expert-validated law articles $A^{\text{exp}}$, flag accepted
2:Updated CLAKG $G^{\prime}=(V^{\prime},E^{\prime},R)$
[…]
4:if $c\notin V^{\prime}$ then add $c$ to $V^{\prime}$
5:end if
6:add / update case-detail nodes for $c$ and link them to $c$
7:obtain $K^{\text{sys}}(c)$ via the LLM-based keyword matching in Section 3.2.2 and expert-validated $K^{\text{exp}}(c)$
8:if accepted = True then $\triangleright$ recommendation accepted
9:  $E^{\prime}\leftarrow E^{\prime}\cup\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{exp}}\}\cup\{(c,k,\texttt{Agree With})\mid k\in K^{\text{exp}}(c)\}$
[…]
11:  $E^{\prime}\leftarrow E^{\prime}\setminus\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{sys}}\}\setminus\{(c,k,\texttt{Agree With})\mid k\in K^{\text{sys}}(c)\}$
12:  $E^{\prime}\leftarrow E^{\prime}\cup\{(c,a,\texttt{Applicable Law})\mid a\in A^{\text{exp}}\}\cup\{(c,k,\texttt{Agree With})\mid k\in K^{\text{exp}}(c)\}$
13:end if
14:return $G^{\prime}=(V^{\prime},E^{\prime},R)$
In practical deployment, the LLM–CLAKG framework operates in a closed-loop manner in collaboration with legal experts, as shown in step 3 of Figure 1 and formalized in Algorithm 2. After the system has produced a recommended law article and a natural-language explanation for a newly emerged case (Section 3.2), experts review the proposed result. This review focuses on both the correctness of the recommended article and the adequacy of the supporting reasoning. In parallel, the system also outputs a small set of key-information nodes (“keywords”) that it deems most relevant to the new case, obtained by reusing the LLM-based matching procedure in Section 3.2.2. Experts inspect and, if necessary, adjust this keyword set as well, so that both the applicable law articles and their supporting key information are explicitly validated.
If the recommendation is confirmed to be appropriate, the corresponding case–law article pair is treated as a new, high-quality supervision signal for the knowledge graph. Concretely, the case node of the current judgment is added to CLAKG, and one or more “Applicable Law” edges are created between this case node and the validated law-article nodes. In addition, the occurrence time, prosecution reason, and a concise description of the case facts stored as case-detail nodes, are linked to the new case node via “Occur in Time”, “Reason” and “Detail” relationship. Building on the expert-validated keyword set, the system then connects the new case to the corresponding key-information nodes by adding “Agree With” edges, again strictly reusing only key-information nodes that already exist in CLAKG. These operations enrich CLAKG with up-to-date case–article and case–keyword links without modifying the underlying LLM or the retrieval algorithm, corresponding to the “accepted” branch in Algorithm 2.
Conversely, if experts deem the recommendation partially or completely incorrect, they can either select the correct articles from the candidate list or manually specify the appropriate law articles. The system then updates CLAKG accordingly: erroneous “Applicable Law” edges can be removed, while new edges are established between the case node and the expert-validated law articles. At the same time, any misleading keyword associations proposed by the system are pruned by removing the corresponding “Agree With” edges, and new “Agree With” edges are created between the case node and the expert-corrected key-information nodes. This correction process is captured by the “corrected” branch in Algorithm 2, where expert feedback directly refines both the case–law article and case–keyword structure of the graph.
This human–machine collaboration enables CLAKG to continuously absorb new cases, updated keyword associations, and law article changes, leading to a denser and more accurate graph. Since subsequent recommendations always start from CLAKG-based retrieval and KG-grounded prompting, an improved graph directly translates into better keyword matching in Section 3.2.2, more precise candidate law-article retrieval in Section 3.2.3, and stronger factual grounding for the LLM in Section 3.2.4. As a result, the closed-loop renewal of CLAKG—covering both law-article and key-information links for new cases—gradually improves recommendation quality and helps mitigate hallucinations in future LLM outputs.
@@ 4.1 Datasets
While the general architecture of our framework is designed to be not limited to a specific jurisdiction in principle, in this study we instantiate and evaluate it on Chinese criminal law as a concrete case study. Specifically, the law-article dataset employed in this section is derived from the Amendment (XI) to the Criminal Law of the People’s Republic of China [39], ratified during the 24th session of the Standing Committee of the 13th National People’s Congress. This amendment is systematically structured into two sections: the General Provisions and the Specific Provisions, encompassing a total of 452 articles—101 in the General law articles and 351 in the Specific law articles. In other words, our law article set already covers the entire universe of criminal law articles under Amendment (XI), and the number of law-article nodes in CLAKG directly reflects the full statutory space relevant to this domain; it cannot be meaningfully “expanded” without moving beyond Chinese criminal law itself. The CLAKG was constructed utilizing the methodology in Section 3.1. We discuss in Section 5 the challenges and potential of adapting this framework to other legal systems (e.g., common-law jurisdictions) and languages.
@@
The criminal case data analyzed in this study were obtained from the China Judgments Online [40] platform. Figure 7 presents one of the cases. In light of the fact that the Amendment (XI) to the Criminal Law came into force on March 1, 2021, all judicial decisions used herein pertain to cases adjudicated after this date. During this period, due to concerns over information security and privacy protection, the number of publicly available criminal case judgments was limited, and the publication policy is outside the control of the authors. As a consequence, certain law articles correspond to only a very small number of cases, while some provisions have no publicly accessible cases at all. To mitigate this constraint, we exhaustively collected all publicly accessible criminal judgments under Amendment (XI) that matched our statutory subset, yielding a dataset of 1,170 case records covering 85 distinct specific-law articles. Among them, 997 case records were designated for the training set and 173 for the test set. The training set encompassed 49 distinct law articles, whereas the test set covered 85 law articles (zero-shot number = 36). Despite this moderate cardinality in absolute terms, the resulting law article recommendation task is already highly non-trivial: the label space is large and severely imbalanced, and, as shown in Section 4.4, several classical text-classification baselines tend to collapse to the majority label. Throughout all experiments, the retriever in our framework still operates over all 452 law-article nodes in CLAKG, so retrieving the correct article is not artificially simplified by a small candidate pool. The Adjudicated Cases Knowledge Graph was subsequently constructed based on the training data, following the approach outlined in Section 3.1.